A solid, pure substance containing just carbon and hydrogen is found to contain 23.6g of carbon and 3.45 g of hydrogen. How many grams of carbon will be in a 6.45 g sample of the solid?
1
Expert's answer
2017-09-12T01:42:52-0400
Solid sample is - carbon (23,6g) + hydrogen (3,45g) = 27,05g (100%)
Carbon content in substance (%) is : 23,6/27,05 = 0,8725 * 100%= 87,25%
In 6,45g solid sample will be : 6,45*0,8725 = 5,628g of carbon